# Rule-based Data Selection for Large Language Models

### v2

@@ 5.1 Experiments Setup
Note that for uniform sampling and the methods involving randomness in the selection of rules, we considered 3 independent trials and averaged the results (see details in Appendix A.6.3).
5.2 General Continued Pre-training
We selected 20K samples from our data source using the methods described above for continued pre-training of Pythia-1B, then benchmarked the model’s performance. The choice of 20K samples was constrained by our GPU resources. Despite 20K being significantly smaller than the pre-training corpus size, we still observed improvements in benchmark results shown in Table 1, with DPP leading in most metrics. We anticipate these differences will be more obvious in domain-specific fine-tuning settings, demonstrated in 5.3 below.
Method ARC-Easy ARC-Challenge Winogrande MMLU SST-2 Average
Pythia-1B 59.8 25.2 53.5 25.6 49.0 42.6
Uniform Sampling 59.4 24.7 53.6 25.6 49.3 42.5
No Rule 59.6 25.1 53.7 25.6 49.2 42.6
DSIR 60.5 25.3 53.3 25.7 49.9 42.9
QuRating Rules 59.8 25.2 54.2 25.8 49.4 42.8
GPT-Uncorrelated 59.6 24.9 53.2 25.8 49.4 42.6
All 50 Rules 60 25.3 53 26.1 49.3 42.7
Random 10 Rules 60 25.1 53.7 25.8 49.5 42.8
DPP 10 Rules 60 25.7 54 26.2 50.1 43.2
Table 1: General continued pre-training of Pythia-1B using 20K selected data samples from SlimPajama (bold text denotes the first place and underlined text denotes the second place). The first row shows the original model’s performance without further training.
5.3 Domain-specifc Fine-tuning
We now focus on domain-specific fine-tuning across four domains: IMDB, Medical, Math, and Code. By selecting 20K domain-related data samples from our source for model training, we aim to enhance domain-specific task performance. As demonstrated in Tables 2 and 3, domain-specific fine-tuning yields more significant improvements than general continued pre-training, where the latter often needs larger datasets to enhance performance due to the broader nature of the training data. Notably, rule-based methods consistently outperform rule-free approaches in general, especially when comparing against Random Select and No Rule. Among all rule-based methods, QuRating underperforms. As previously noted, such human-designed rules are inherently limited due to varying preferences of designers, introducing bias in rules. Additionally, human designers may not capture the data-dependent correlation between rules effectively. The GPT-Uncorrelated rules face a similar issue where the rule selection process is entirely independent of the data. In contrast, our framework begins by automatically generating a diverse set of rules and then selecting an orthogonal subset. Furthermore, our method employs the data-dependent score vector to represent rules and utilizes a quantitative measure to accurately assess their correlation.
Within our framework, DPP demonstrated superior performance compared to using all 50 rules or selecting 10 rules randomly. This aligns with our argument of the importance of rule orthogonality, as well as the intuition that the optimal $r$ is not near the boundaries (both validated by the previous experiments on Section 4). This underscores the effectiveness of a rule-based strategy, which introduces more balanced diversity in the data rating aspects and selects better training data. Furthermore, it also demonstrates that our application of DPP in rule selection effectively identifies a core set of high-quality rules, thereby enhancing data quality and ultimately improving model performance.
Method IMDB Medical
Pythia-1B 44.5 21.4 34.2 23.0 26.2
Uniform Sampling 43.9 23.0 42.1 22.5 28.9
No Rule 51.1 23.1 42.6 22.0 29.2
DSIR 50.2 22.5 32.4 17.0 23.9
QuRating 47.7 21.3 42.2 22 28.5
GPT-Uncorrelated 50.9 23.7 42 22.7 29.4
All 50 Rules 51 23.1 42.6 23 29.6
Random 10 Rules 51.7 24 41.2 23.5 29.6
DPP 10 Rules 53.5 24.6 43.3 26.8 31.6
@@ 5.3 Domain-specifc Fine-tuning
Table 2: IMDB & Medical fine-tuning on Pythia-1B, each using 20K selected data samples from SlimPajama. The first row shows the original model’s performance without further training.
Method Math Code
Llama3-8B 41 39.6 34 38.2 46.3 42.9 44 48.4 45.4
Uniform Sampling 40.5 39.2 35 38.2 38.7 38.2 38.2 39.7 38.7
No Rule 42.3 37.4 37 38.9 45.1 43.9 42.8 52.1 45.9
DSIR 41.5 41.1 34 38.9 45.1 43.6 49.1 52.2 47.5
QuRating 41.5 38.1 35 38.2 43.2 43.4 40.5 45.6 43.1
GPT-Uncorrelated 41.4 39 37.3 39.2 41.2 43.5 39.6 48.6 43.2
All 50 Rules 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.8 35.2 39.3 48.5 41 46.6 48.1 46
DPP 10 Rules 43.7 40.6 38 40.8 50.5 44.2 46.9 52.7 48.6
Table 3: Math & Code fine-tuning on Llama3-8B, each using 20K selected data samples from SlimPajama. The first row shows the original model’s performance without further training.
6 Conclusion
We have introduced an automated, rule-based framework for selecting high-quality LLM data, utilizing LLMs to generate a diverse set of rules and the DPP method to eliminate redundancy. Our work is the first to introduce an automated rule evaluation metric and we also propose a rule-based selection pipeline that demonstrates substantial generalizability across various settings, effectively overcoming the limitations of human-designed rules and addressing the challenges associated with the lack of robust rule evaluations. We first demonstrated that our approach enhances the accuracy of data ratings using a dataset with given ground truth scores. Then we conduct experiments that train LLMs with selected data and have shown that our method outperforms various other approaches, both in general pre-training and fine-tuning across four domains. The results indicate that our method successfully generates high-quality, diverse rules, and thereby improves quality of selected data, which in turn leads to improved model performance after trained with the chosen data.
Appendix A Appendix
A.1 Orthogonality Measures
@@
where ${\bm{v}}_{i}$ are the columns of $\bar{{\bm{S}}}$. The determinant of $\bar{{\bm{S}}}^{\top}\bar{{\bm{S}}}$ geometrically represents the squared volume of the parallelepiped formed by the columns of $\bar{{\bm{S}}}$ (Kulesza et al., 2012). We normalize by the product of the vector norms since both the magnitude of the vectors and their mutual correlation influence the volume: larger norms increase the volume, whereas higher correlation reduces it. Thus, after normalization, the value of Vol serves as an indicator of the overall orthogonality among the column vectors of $\bar{{\bm{S}}}$. The phenomena under the usage of this measure are similar to the ones under 1. Therefore we only presented results using the rule correlation.
A.2 DPP Sampling
@@
DPP Sampling Algorithm: The sampling algorithm can be found in Algorithm 1 of Kulesza et al. (2012). The sampling process starts by decomposing the kernel matrix ${\bm{K}}$ and involves two main stages: 1. Selecting eigenvectors by sampling from a Bernoulli distribution based on the eigenvalues, and 2. Sampling a subset from the ground set using an iterative conditional distribution method to ensure diversity, as detailed in (Kulesza et al., 2012). We utilize the DPPy Python library (Gautier et al., 2019) for efficient DPP initialization and sampling.
Time Complexity: Finding the submatrix (subset of columns) of a matrix to maximize the orthogonality is NP-hard (Civril & Magdon-Ismail, 2007; Kulesza et al., 2012). DPP provides us a relatively good solution. In practice, the computational complexity of sampling from a DPP depends primarily on the eigendecomposition of the kernel matrix ${\bm{K}}$. In our case, ${\bm{K}}\in\mathbb{R}^{R\times R}$ and therefore it requires $O(R^{3})$ time, where $R$ is the number of rules. In the DPPy package (Gautier et al., 2019) it uses the spectral sampler by default, so the actual run-time of our DPP implementation is $O(R^{3})$.
DPP Sampling for Data Selection: We noticed that in a concurrent work Yang et al. (2024), the authors also use DPP to perform data selection, but directly applied to the data itself. However, the approach to directly perform data selection using DPP requires the computation based on the kernel matrix with dimension $N$ (number of samples), which is usually huge in the context of LLM data. Moreover, while DPP inherently prioritizes diversity in data selection, it does not address other quality dimensions. In contrast, our rule-based approach assesses multiple aspects of data quality, ensuring a more comprehensive and robust selection process.
A.3 Stochastic data selection: Gumbel top-$k$ trick:
Imagine the cases where the target dataset distribution shows a long-tail pattern with respect to our quality measure, using a deterministic quality score as the cutoff could exclude many possibly valuable data (Albalak et al., 2024). Hence, our stochastic sampling in 3 effectively balances the quality and diversity of the selected data. Nonetheless, instead of doing actual sampling according to Equation 3, we use the Gumbel top-$k$ trick similar as in (Wettig et al., 2024), which is a sampling technique used to efficiently and probabilistically select the top-$k$ items from a discrete probability distribution. Specifically, each item $i$ in the distribution is assigned a score using the formula:
@@ A.3 Stochastic data selection: Gumbel top-$k$ trick:
where $p_{i}$ is the probability of item $i$, and $g_{i}$ is a noise term drawn from a Gumbel distribution, which can be generated using $g_{i}=-\log(-\log(u_{i}))$. In other words, we could add a Gumbel noise vector to the log of the sampling probability in Equation 3 and then choose the top-$k$ data points with the highest sums. This is statistically equivalent to sampling according to Equation 3 (Kool et al., 2019).
A.4 Limitations and Future Directions
Adjusting hyperparameters. Recall that our hyperparameter $r$ determines the number of rules selected for rating, influencing the diversity and coverage of the selected rules. We have explored the effect of $r$ in Section 4 and also in Appendix A.6.5. We leave a comprehensive study of its optimal values for future work.
Data sampling method. There are variations of the stochastic top-$k$ sampling, such as incorporating a temperature parameter $\tau$ (see Wettig et al. (2024)). Replacing equation 3 with its variations or exploring other data sampling methods represents another research direction.
Other rule evaluations metrics. We have proposed multiple metrics in 1 and A.1 to measure rule quality, but all based on the correlation/orthogonality of rules. Evaluating rules from other aspects is another intriguing topic for future work.
A.5 Appendix for Evaluation A
A.5.1 Bradley Terry Model
The Bradley-Terry model is a probabilistic model used to estimate the latent “strength” of teams based on pairwise competitions. The model is parameterized as follows:
$\mathbb{P}(i\text{ beats }j)=\frac{v_{i}}{v_{i}+v_{j}}=\frac{e^{\beta_{i}}}{e^{\beta_{i}}+e^{\beta_{j}}},$ (6)
where exponential functions are used to model the scores $v_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}e^{\beta_{i}}$ and $v_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}e^{\beta_{j}}$. In other words, the difference of their scores determines the the log-odds of team $i$ beating team $j$. Sometimes an intercept term $\alpha$ is added to adjust for any influence of the order (for example, imagine that $i$ is the home team and has home-court advantage), then the probability becomes
$\mathbb{P}(i\text{ beats }j)=\frac{e^{\alpha+\beta_{i}}}{e^{\alpha+\beta_{i}}+e^{\beta_{j}}},$ (7)
The most straightforward method for estimating these parameters is through maximum likelihood estimation, which optimizes the likelihood of the observed outcomes based on the model and its parameters. More details can be found in Bradley & Terry (1952); Hunter (2004).
A.5.2 Error Metrics
Ranking-difference error. To assess the deviation of rating scores from the ground truth, instead of using the mean squared error in 4, an alternative intuitive approach is to compare the rankings derived from the data scores with those of the ground truth. This approach is based on the premise that for data selection purposes, if two sets of scores yield identical rankings, they will select the same high-scoring data samples. An example of such a ranking metric is the Kendall rank correlation coefficient (Kendall’s tau) (Kendall, 1938). However, we opted against this type of metric for two critical reasons: First, it lacks the granularity needed to evaluate errors effectively. For instance, two sets of scores like [0.01, 0.98, 0.99] and [0.01, 0.02, 0.03] share exactly the same ranking yet differ significantly in their actual scores. Second, our method involves stochastic data selection, not a straightforward top-$k$ selection, meaning that a higher score increases the likelihood of a data point being chosen. Hence, a ranking difference, which overlooks the absolute values of scores and focuses solely on their relative comparisons, is not ideal here.
A.5.3 Rating scheme Variation: Individual rating
Here we present the results after replacing the pair-wise rating with the direct individual rating in Section 4:
Figure 4: (a) and (b): Pearson correlation of the rule correlation $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$, for IMDB and CommonCrawl datasets respectively. (c) and (d): Distribution of MSE from $10^{6}$ possible rule subsets with size $r$, for IMDB and CommonCrawl datasets respectively. Two vertical lines represent the MSE values of QuRating and NoRule.
Figure 5: (a) Winning rate of DPP-selected rules compared to QuRating’s four rules and the NoRule setting, based on MSE across 100 DPP trials. (b) Comparison of DPP rule correlation vs. random rule correlation (averaged over 100 trials). (c) Comparison of MSE between DPP-selected and randomly selected rules, averaged across 100 trials. Plots (a), (b), and (c) display results for the IMDB dataset, while (d), (e), and (f) for the CommonCrawl dataset.
A.5.4 Rater Model Size Variation: Llama3-70B-Instruct
Here we present the results after replacing the rater model from Llama3-8B-Instruct model with the stronger Llama3-70B-Instruct in Section 4:
Figure 6: (a): Pearson correlation of the rule correlation $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$ (b): Distribution of MSE from $10^{6}$ possible rule subsets with size $r$. Two vertical lines represent the MSE values of QuRating and NoRule.
Figure 7: (a) Winning rate of DPP-selected rules compared to QuRating’s four rules and the NoRule setting, based on MSE across 100 DPP trials. (b) Comparison of rule correlation between DPP-selected and randomly selected rules, averaged across 100 trials. (c) Comparison of MSE between DPP-selected and randomly selected rules, averaged across 100 trials.
A.5.5 Rule Generator Variation: Claude-3.5-Sonnet
@@ A.5.4 Rater Model Size Variation: Llama3-70B-Instruct
To verify that GPT-4 is a reliable rule generator, we compare it with Claude-3.5-Sonnet. For each of the five tasks (General, IMDB, Medical, Math, Code), we prompt GPT and Claude to generate 100 rules for each, and then study the distribution of the rules. Specifically, we use Sentence-Transformer (Reimers, 2019) to generate the embedding vectors and then use PCA to project them onto the top two principal components for 2-dimensional visualization. From Figure 8 below, we observe that the two groups of rules generated by the two models completely overlap, demonstrating no distinct separation. This suggests that GPT-4 functions effectively as an unbiased rule generator.
(a) General Rules
(b) IMDB Rules
(c) Medical Rules
(d) Math Rules
(e) Code Rules
@@
Now we have compared the rule generators using GPT and Claude above. In order to address potential biases in rule generation by these large language models (LLMs) compared to human-generated rules, we prompt GPT-4 to generate 133 ethical and safety rules and compare the GPT-generated rules with the public and standard constitutions in Huang et al. (2024) (we make the rules all start from “Choose the response that” for a fair comparison). We asked 3 authors who have not seen the constitutions in Huang et al. (2024) to distinguish the rules blindly. We get an average accuracy of $20.7\%$, suggesting it is indeed hard to distinguish between rules generated by GPT and those designed by humans. All these discussions underscores the potential of GPT as a reliable rule generator that is capable of producing rules that are comparable to those crafted by human experts.
A.5.6 Prompts and Generated Rules

### v3

@@ Appendix B Ablation Study
Note that for uniform sampling and the methods involving randomness in the selection of rules, we considered 3 independent trials and averaged the results.
Method IMDB Medical Math Code
SA
accuracy
college
medicine
professional
medicine
medical
genetics
Medical
Average
elementary
school
high
school
college
math
Math
Average
human
-eval
mbpp
multiple
-py
multiple
-cpp
Code
Average
GPT-selected 10 Rules 51.6 21.9 42.2 24 29.4 42.5 40.3 36 39.6 40.8 43.4 42.8 50.3 44.3
All 50 Rules 51 23.1 42.6 23 29.6 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 51.7 24 41.2 23.5 29.6 42.9 39.8 35.2 39.3 48.5 41 46.6 48.1 46
DPP 10 Rules 53.5 24.6 43.3 26.8 31.6 43.7 40.6 38 40.8 50.5 44.2 46.9 52.7 48.6
@@
Table 2: Fine-tuning Llama3-8B on IMDB, Medical, Math and Code domains, each using 20K selected data samples from SlimPajama. The first row shows the performance of the original backbone model (Pythia-1B for IMDB and Medical and Llama3-8B for Math and Code).
Rule Correlation of selected rules. Here we provide in Table 3 the rule indices (in the range $\{0,1,\dots,49\}$) for the rules selected by DPP and random selection (in one trial). The complete list of all generated $50$ rules for each domain is provided in I.8. For each set of selected rules, we also calculate their rule correlation value $\rho$ (defined in 1). We confirm that indeed DPP selects rules with lower rule correlation than random selected rules.
Table 3: Rule correlation and indices of the selected rules by DPP and random selection (two trials “A” and “B”).
Domain Method Rule Indices
IMDB DPP 10 rules ($\rho=0.42$) [2, 3, 13, 21, 28, 36, 37, 45, 46, 49]
Random 10 Rules A ($\rho=0.53$) [2, 6, 10, 11, 15, 21, 28, 42, 43, 48]
Random 10 Rules B ($\rho=0.51$) [1, 9, 12, 14, 25, 26, 27, 37, 38, 40]
Medical DPP 10 rules ($\rho=0.55$) [1, 9, 10, 25, 29, 30, 32, 38, 42, 47]
Random 10 Rules A ($\rho=0.69$) [11, 13, 14, 16, 25, 33, 34, 43, 45, 49]
Random 10 Rules B ($\rho=0.66$) [6, 17, 20, 28, 29, 37, 40, 41, 47, 48]
Math DPP 10 rules ($\rho=0.40$) [0, 4, 13, 26, 27, 31, 33, 38, 44, 45]
Random 10 Rules A ($\rho=0.65$) [0, 2, 11, 16, 17, 18, 27, 28, 34, 39]
Random 10 Rules B ($\rho=0.61$) [3, 4, 25, 20, 23, 13, 15, 24, 35, 39]
Code DPP 10 rules ($\rho=0.54$) [2, 3, 13, 21, 28, 36, 37, 45, 46, 49]
Random 10 Rules A ($\rho=0.59$) [5, 7, 10, 13, 17, 19, 21, 26, 30, 34]
Random 10 Rules B ($\rho=0.58$) [2, 4, 8, 14, 16, 20, 23, 33, 37, 44]
In Table 4 below, we calculate the rule correlation selected by GPT. By comparing to the rule correlations of DPP-selected rules in Table 3, we see that although we prompt GPT-4 to select “uncorrelated” rules, the rule correlations of the selected 10 rules are still higher than our DPP-selected rules.
Table 4: Rules selected by GPT and their rule correlation values.
Domain Method Rule Indices
IMDB GPT selected 10 rules ($\rho=0.67$) [0, 1, 4, 10, 13, 17, 25, 31, 40, 49]
Medical GPT selected 10 rules ($\rho=0.40$) [0, 4, 7, 11, 15, 24, 29, 34, 42, 49]
Math GPT selected 10 rules ($\rho=0.56$) [0, 3, 7, 11, 17, 24, 28, 38, 44, 48]
Code GPT selected 10 rules ($\rho=0.65$) [0, 4, 9, 12, 16, 23, 29, 34, 43, 49]
Appendix C Limitations and Future Directions
Adjusting hyperparameters. Recall that our hyperparameter $r$ determines the number of rules selected for rating, influencing the diversity and coverage of the selected rules. We have explored the effect of $r$ in Section 4 and also in Appendix I.5. We leave a comprehensive study of its optimal values for future work.
Data sampling method. There are variations of the stochastic top-$k$ sampling, such as incorporating a temperature parameter $\tau$ (see wettig2024qurating). Replacing (4) with its variations or exploring other data sampling methods represents another research direction.
Other rule evaluations metrics. We propose multiple metrics (in Equations (1) and (11) below) to measure rule quality, but all based on the correlation/orthogonality of rules. Evaluating rules from other aspects is another intriguing topic for future work.
Appendix D Orthogonality Measures
@@
where ${\bm{v}}_{i}$ are the columns of $\bar{{\bm{S}}}$. The determinant of $\bar{{\bm{S}}}^{\top}\bar{{\bm{S}}}$ geometrically represents the squared volume of the parallelepiped formed by the columns of $\bar{{\bm{S}}}$ (kulesza2012determinantal). We normalize by the product of the vector norms since both the magnitude of the vectors and their mutual correlation influence the volume: larger norms increase the volume, whereas higher correlation reduces it. Thus, after normalization, the value of Vol serves as an indicator of the overall orthogonality among the column vectors of $\bar{{\bm{S}}}$. The phenomena under the usage of this measure are similar to the ones under 1. Therefore we only presented results using the rule correlation $\rho$.
Appendix E DPP Sampling
@@
DPP Sampling Algorithm: The sampling algorithm can be found in Algorithm 1 of kulesza2012determinantal. The sampling process starts by decomposing the kernel matrix $\bm{K}$} and involves two main stages: 1. Selecting eigenvectors by sampling from a Bernoulli distribution based on the eigenvalues, and 2. Sampling a subset from the ground set using an iterative conditional distribution method to ensure diversity, as detailed in (kulesza2012determinantal). We utilize the DPPy Python library (GPBV19) for efficient DPP initialization and sampling.
Time Complexity: Finding the submatrix (subset of columns) of a matrix to maximize the orthogonality is NP-hard (civril2007finding; kulesza2012determinantal). DPP provides us a relatively good solution. In practice, the computational complexity of sampling from a DPP depends primarily on the eigendecomposition of the kernel matrix $\bm{K}$}. In our case, ${\bm{K}}\in\mathbb{R}^{R\times R}$ and therefore it requires $O(R^{3})$ time, where $R$ is the number of rules. In the DPPy package (GPBV19) it uses the spectral sampler by default, so the actual run-time of our DPP implementation is $O(R^{3})$.
DPP Sampling for Data Selection: We noticed that in a concurrent work (yang2024p3), the authors also use DPP to perform data selection, but directly applied to the data itself. However, the approach to directly perform data selection using DPP requires the computation based on the kernel matrix with dimension $N$ (number of samples), which is usually huge in the context of LLM data. Moreover, while DPP inherently prioritizes diversity in data selection, it does not address other quality dimensions. In contrast, our rule-based approach assesses multiple aspects of data quality, ensuring a more comprehensive and robust selection process.
Appendix F Stochastic data selection: Gumbel top-$k$ trick:
Imagine the cases where the target dataset distribution shows a long-tail pattern with respect to our quality measure, using a deterministic quality score as the cutoff could exclude many possibly valuable data (albalak2024survey). Hence, our stochastic sampling in Equation (4) effectively balances the quality and diversity of the selected data. Nonetheless, instead of doing actual sampling according to Equation (4), we use the Gumbel top-$k$ trick similar as in (wettig2024qurating), which is a sampling technique used to efficiently and probabilistically select the top-$k$ items from a discrete probability distribution. Specifically, each item $i$ in the distribution is assigned a score using the formula:
@@ Appendix F Stochastic data selection: Gumbel top-$k$ trick:
where $p_{i}$ is the probability of item $i$, and $g_{i}$ is a noise term drawn from a Gumbel distribution, which can be generated using $g_{i}=-\log(-\log(u_{i}))$. In other words, we could add a Gumbel noise vector to the log of the sampling probability in (4) and then choose the top-$k$ data points with the highest sums. This is statistically equivalent to sampling according to (4) (kool2019stochastic).
Appendix G Rule Generator Variation: Claude-3.5-Sonnet
@@
To verify that GPT-4 is a reliable rule generator, we compare it with Claude-3.5-Sonnet. For each of the five tasks (General, IMDB, Medical, Math, Code), we prompt GPT and Claude to generate 100 rules for each, and then study the distribution of the rules. Specifically, we use Sentence-Transformer (reimers2019sentence) to generate the embedding vectors and then use PCA to project them onto the top two principal components for 2-dimensional visualization. From Figure 4 below, we observe that the two groups of rules generated by the two models completely overlap, demonstrating no distinct separation. This suggests that using strong LLMs such as GPT or Claude does not introduce clear model-specific bias during rule generation.
(a) IMDB Rules
(b) Medical Rules
(c) Math Rules
(d) Code Rules
@@
We have compared the rule generators using GPT and Claude above, both are LLMs. In order to address potential biases in rule generation by these LLMs compared to human-generated rules, we prompt GPT-4 to generate 133 ethical and safety rules and compare the GPT-generated rules with the public and standard constitutions in (huang2024collective) (we make the rules all start from “Choose the response that” for a fair comparison). We asked 3 authors who have not seen the constitutions in (huang2024collective) to distinguish the rules blindly. We get an average accuracy of $20.7\%$, suggesting it is indeed hard to distinguish between rules generated by GPT and those designed by humans. All these discussions underscore the potential of GPT as a reliable rule generator that is capable of producing rules that are comparable to those crafted by human experts.
Appendix H Appendix for Preliminary Experiments
H.1 Error Metrics
Ranking-difference error. To assess the deviation of rating scores from the ground truth, instead of using the mean squared error in 5, an alternative intuitive approach is to compare the rankings derived from the data scores with those of the ground truth. This approach is based on the premise that for data selection purposes, if two sets of scores yield identical rankings, they will select the same high-scoring data samples. An example of such a ranking metric is the Kendall rank correlation coefficient (Kendall’s tau) (kendall1938new). However, we opted against this type of metric for two critical reasons: First, it lacks the granularity needed to evaluate errors effectively. For instance, two sets of scores like [0.01, 0.98, 0.99] and [0.01, 0.02, 0.03] share exactly the same ranking yet differ significantly in their actual scores. Second, our method involves stochastic data selection, not a straightforward top-$k$ selection, meaning that a higher score increases the likelihood of a data point being chosen. Hence, a ranking difference, which overlooks the absolute values of scores and focuses solely on their relative comparisons, is not ideal here.
H.2 Other domains: Medical, Math and Code
In our preliminary experiments (Section 4), we focused on the IMDB dataset to illustrate the effectiveness of our method for rule-based rating. Here, we extend the same Evaluation A analysis to three additional domains: Medical, Math, and Code. The corresponding results are shown in Figure 5. These include the correlation between rule diversity and rating accuracy, comparisons between rule-based and rule-free settings, and evaluations of DPP-selected rules versus human-designed and randomly selected ones. The findings further confirm the generalizability of our method across diverse domains.
Figure 5: Each row corresponds to a domain: (top) Medical, (middle) Math, and (bottom) Code. (a): Pearson correlation between rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$. (b): Distribution of MSE from $10^{6}$ possible rule subsets with size $r$, where vertical lines represent QuRating and NoRule. (c): Winning rate of DPP-selected rule subsets compared to QuRating and NoRule baselines based on MSE. (d,e): Comparison of DPP-selected versus randomly selected rules on rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$, respectively.
H.3 General Dataset: CommonCrawl
In Figure 6, we observe that QuRating rules here have better performance than on IMDB dataset: it can at least surpass most of the randomly selected rules. This is reasonable since the QuRating themselves are general high-level rules. It suggests that while QuRating rules are effective for general pre-training data, they lack the flexibility to adapt to other settings or domains. When compared to DPP, interestingly, DPP underperforms QuRating when $r$ is too small or too large.
Figure 6: Evaluation on the CommonCrawl dataset. (a): Pearson correlation between rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$. (b): Distribution of MSE across $10^{6}$ random rule subsets with size $r$, with vertical lines indicating MSE values of QuRating and NoRule. (c): Winning rate of DPP-selected rule subsets compared to QuRating and NoRule. (d,e): Comparison of DPP-selected versus randomly selected rules on rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$, averaged over 100 trials.
H.4 Empirical Distribution of $\rho$, $\epsilon$, and Ground Truth Ratings
To verify that the rule correlation $\rho(\bar{{\bm{S}}})$, the MSE $\epsilon(\bar{{\bm{S}}})$, and the ground truth ratings $S_{GT}$ are not trivially concentrated around a narrow range or even single value, we plot their empirical distributions across different domains. As shown in Figure 7, each metric exhibits substantial variance, indicating meaningful diversity among rule subsets and human-provided quality scores. This supports the validity of using these signals for evaluating rule selection strategies.
Figure 7: Distributions of rule correlation $\rho(\bar{{\bm{S}}})$, MSE $\epsilon(\bar{{\bm{S}}})$, and ground truth scores across four datasets: IMDB, Medical, Math, and Code.
H.5 Prompts and Generated Rules